# Locating Superspreaders in SI Cascades

### 05 Mar 2024 v1

@@ 5.1 First derivative is approximately the size of the cut
Since $G$ is a tree, we have for each $v\in\partial{\mathcal{I}}(t)$ that $v$ has exactly one infected neighbor at time $t$. Hence $T(v)-t\sim\mathrm{Exp}(\lambda)$, so
$\operatorname{\mathbb{P}}(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t})=\operatorname{\mathbb{P}}(T(v)-t\in[0,n^{-\beta}]|{\mathcal{F}}_{t})=1-e^{-\lambda/n^{\beta}}.$
By the conditional independence of infection times in $\partial{\mathcal{I}}(t)$ with respect to ${\mathcal{F}}_{t}$, we have that $|{\mathcal{I}}_{1}|\sim\mathrm{Bin}(|\partial{\mathcal{I}}(t)|,1-e^{-\lambda/n^{\beta}})$. We can then bound
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-|\partial{\mathcal{I}}(t)|\left(1-e^{-\lambda/n^{\beta}}\right)\right|\geq\frac{1}{2}n^{\gamma}\right|{\mathcal{F}}_{t}\right)$ $\stackrel{{\scriptstyle(a)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{|\partial{\mathcal{I}}(t)|\left(1-e^{-\lambda/n^{\beta}}\right)+n^{\gamma}/6}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{\lambda n^{1-\beta}+n^{\gamma}/6}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}o\left(\frac{1}{n}\right).$ (5.1)
Above, $(a)$ is due to Bernstein’s inequality, $(b)$ uses the upper bounds $|\partial{\mathcal{I}}(t)|\leq n$ and $1-e^{-\lambda/n^{\beta}}\leq\lambda/n^{\beta}$, and $(c)$ follows since $\gamma>(1-\beta)/2$.
To prove the claim, it remains to study $|\partial{\mathcal{I}}(t)|(1-e^{-\lambda/n^{\beta}})$. We have that
$\left||\partial{\mathcal{I}}(t)|\left(1-e^{-\lambda/n^{\beta}}\right)-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\stackrel{{\scriptstyle(d)}}{{=}}\left|1-e^{-\lambda/n^{\beta}}-\frac{\lambda}{n^{\beta}}\right||\partial{\mathcal{I}}(t)|\stackrel{{\scriptstyle(e)}}{{\leq}}\lambda^{2}n^{1-2\beta}\stackrel{{\scriptstyle(f)}}{{\leq}}\frac{1}{2}n^{\gamma},$ (5.2)
where $(d)$ follows since $|\partial{\mathcal{I}}(t)|=|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|$ in the case where $G$ is a tree [ref this fact], $(e)$ is due to a Taylor approximation and since $|\partial{\mathcal{I}}(t)|\leq n$, and $(f)$ holds for $n$ sufficiently large since $\gamma>1-2\beta$. Together, (5.1) and (5.2) prove the lemma. ∎
The same ideas can be used to characterize $|{\mathcal{I}}_{1}|$ in general graphs (i.e., non-trees) as well. The key difference in the general setting is that vertices in $\partial{\mathcal{I}}(t)$ may have multiple infected neighbors. For high-degree vertices, this effect is most pronounced; indeed, as Lemma [ref] shows, it may be the case that a high-degree $v\in\partial{\mathcal{I}}(t)$ has up to $\sqrt{n}$ infected neighbors. Accounting for these qualitative differences from the tree case leads to slight differences in our results for general graphs, as the following result shows.
Lemma 5.2.
Let $G$ be a graph on $n$ vertices, let $\beta>1/2$ and assume that $\gamma>\max\{(1-\beta)/2,1-2\beta\}$. Recall the event ${\mathcal{A}}$ from Definition 4.5. Then for any $t\geq 0$, it holds that
($\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)=o\left(\frac{1}{n}\right).$
Proof.
As in the proof of Lemma [ref], define the event
${\mathcal{A}}_{t}:=\left\{\text{For all $v\in V\setminus{\mathcal{I}}[0,t]$, it holds that $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq n^{1/2+o(1)}$}\right\}.$
@@
($\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)$ ($\leq\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}_{t}\right|{\mathcal{F}}_{t}\right)$
$=\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)\mathbf{1}({\mathcal{A}}_{t}).$ (5.3)
We therefore proceed by bounding the probability on the right hand side, using the same proof ideas as Lemma 5.1. For each $v\in\partial{\mathcal{I}}(t)$, notice that
$\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v)\sim\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|),$
@@
$\operatorname{\mathbb{P}}\left(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t}\right)=1-\exp\left(-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right).$
Moreover, by the independence of edge weights, we have that $|{\mathcal{I}}_{1}|\stackrel{{\scriptstyle d}}{{=}}\sum_{v\in\partial{\mathcal{I}}(t)}X(v)$, where the $X(v)$’s are independent with $X(v)\sim\operatorname{Bern}\left(1-e^{-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|}\right)$. We can then bound
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}||{\mathcal{F}}_{t}]\right|\geq\frac{1}{2}n^{\gamma}\right|{\mathcal{F}}_{t}\right)$ $\stackrel{{\scriptstyle(a)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-\exp(-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|)\right)+n^{\gamma}/6}\right)$
@@
Above, $(a)$ is due to Bernstein’s inequality. Step $(b)$ uses the inequalities $|\partial{\mathcal{I}}(t)|\leq n$ as well as
$\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-e^{-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|}\right)\leq\frac{\lambda}{n^{\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq\frac{\lambda}{n^{\beta}}\left(dn+nm\right)=n^{1-\beta+o(1)},$ (5.5)
where the second inequality above holds since there are at most $n$ vertices $v$ with $|{\mathcal{N}}(v)|\leq d$ and $m$ vertices $v$ with $|{\mathcal{N}}(v)|\leq n$; the final expression uses that $d=n^{o(1)}$ and $m=n^{o(1)}$. Finally, $(c)$ is due to our assumption that $\gamma>1/2-\beta/2$.
It remains to study $\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}|{\mathcal{F}}_{t}]$. We have that
$\left|\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}||{\mathcal{F}}_{t}]-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|$ $=\left|\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-\exp(\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|)\right)-\frac{\lambda}{n^{\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right|$
$\leq\sum_{v\in\partial{\mathcal{I}}(t)}\left|1-\exp(\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|-\frac{\lambda}{n^{\beta}}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right|$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{\lambda^{2}}{n^{2\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|$
@@
Above, $(c)$ is due to Taylor’s theorem (in particular, the inequality holds as long as $\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|=o(1)$, which is the case since ${\mathcal{A}}_{t}$ holds and $\beta>1/2$). Step $(d)$ uses the bound $\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq n^{1+o(1)}$, which was proved in (5.5). Finally, $(e)$ follows since $\gamma>1-2\beta$.
Together, (5.3), (5.4) and (5.6) prove the lemma. ∎
[next result: bounding $|{\mathcal{I}}_{2}|$ for both trees and general graphs]
Lemma 5.3.
Let $\beta,\gamma>0$ and let $t\geq 0$. If $G$ is a tree, then
$\operatorname{\mathbb{P}}\left(\left.|{\mathcal{I}}_{2}|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)\leq n^{-\min\{\beta,\gamma+2\beta-1\}+o(1)}.$
For general $G$, it holds that
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)\leq n^{-\min\{\beta-1/2,\gamma+2\beta-1\}+o(1)}.$
Proof.
As a shorthand let $\delta:=n^{-\beta}$, and define the event
${\mathcal{B}}_{t}:=\{\mathsf{highdeg}(G)\cap{\mathcal{I}}(t,t+\delta]=\emptyset\}.$
Observe that for each $u\in{\mathcal{I}}_{2}$, there must exist a distinct path $P$ from ${\mathcal{I}}[0,t]$ to $u$ of length at least 2 and weight at most $\delta$ such that only the first vertex in $P$ is an element of ${\mathcal{I}}[0,t]$. Moreover, if ${\mathcal{B}}_{t}$ holds, then all vertices in $P$ must have degree at most $d$. Otherwise, there would exist a path of weight at most $\delta$ connecting a vertex in ${\mathcal{I}}[0,t]$ to some $u\in\mathsf{highdeg}(G)$, which would imply that $T(u)\in(t,t+\delta]$. However, this cannot happen on the event ${\mathcal{B}}_{t}$.
To see the implications of the discussion above, let ${\mathcal{P}}$ be the set of paths $P$ in $G$ such that $|P|\geq 2$, the first vertex of $P$ is the only one in ${\mathcal{I}}[0,t]$, and all vertices in $P$ have degree at most $d$. Since there exists a distinct path $P\in{\mathcal{P}}$ with weight at most $\delta$ for each $u\in{\mathcal{I}}_{2}$, it holds on the event ${\mathcal{B}}_{t}$ that
$|{\mathcal{I}}_{2}|\leq|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}|.$
We proceed by bounding the expectation of the right hand side as follows.
$\operatorname{\mathbb{E}}\left[|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}||{\mathcal{F}}_{t}\right]=\sum_{\ell\geq 2}\sum_{P\in{\mathcal{P}}:|P|=\ell}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})\leq\sum_{\ell\geq 2}|\{P\in{\mathcal{P}}:|P|=\ell\}|(\lambda\delta)^{\ell}$ (5.7)
Above, the final inequality holds since $\mathsf{weight}(P)\sim\mathrm{Gamma}(|P|,\lambda)$. We continue by bounding the number of paths in ${\mathcal{P}}$ of length $\ell$ through elementary combinatorial arguments. If $v\in{\mathcal{I}}[0,t]$ has degree at most $d$, the number of paths $P\in{\mathcal{P}}$ which begin at $v$ and have length $\ell$ is $d^{\ell}$. As there are at most $n$ vertices in ${\mathcal{I}}[0,t]$ of degree at most $d$, the total number of $P\in{\mathcal{P}}$ of length $\ell$ starting at such a vertex is at most $nd^{\ell}$. On the other hand, if $v\in{\mathcal{I}}[0,t]$ is in $\mathsf{highdeg}(G)$, then the total number of paths of length $\ell$ beginning at $v$ is at most $nd^{\ell-1}$, since the degree of $v$ is at most $n$ but all other vertices in the path have degree at most $d$. As there are $m$ vertices in $\mathsf{highdeg}(G)$, the number of such paths is at most $mnd^{\ell-1}$. Combining both cases, we have that
|$|\{P\in{\mathcal{P}}:|P|=\ell\}|\leq nd^{\ell}+mnd^{\ell-1}\leq n^{1+o(1)}d^{\ell}.$
Substituting this bound into (5.7) shows that
$\operatorname{\mathbb{E}}\left[|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}||{\mathcal{F}}_{t}\right]\leq\sum_{\ell\geq 2}n^{1+o(1)}d^{\ell}(\lambda\delta)^{\ell}=n^{1-2\beta+o(1)},$
where the final expression follows from the geometric sum formula and since $d\lambda\delta=n^{-\beta+o(1)}$. Putting everything together, we have that
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)\leq\operatorname{\mathbb{P}}\left(|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}|\geq n^{\gamma}|{\mathcal{F}}_{t}\right)\leq n^{-(\gamma+2\beta-1)+o(1)},$ (5.8)
where the final bound is due to Markov’s inequality. To prove the claim, it remains to study the probability of ${\mathcal{B}}_{t}$. If $G$ is a tree, then
$\operatorname{\mathbb{P}}({\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})\leq\sum_{v\in\mathsf{highdeg}(G)\setminus{\mathcal{I}}[0,t]}\operatorname{\mathbb{P}}\left(T(v)\in(t,t+\delta]|{\mathcal{F}}_{t}\right)\leq m\lambda\delta=n^{-\beta+o(1)}.$ (5.9)
Above, the second inequality holds since there are $m$ high-degree vertices and since $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda)$ by Lemma [ref]. Together, (5.8) and (5.9) prove the lemma in the case where $G$ is a tree.
Next, suppose that $G$ is not necessarily a tree. Recalling the event ${\mathcal{A}}$ from Definition 4.5, we have by Lemma [ref] that
We can therefore bound
($\operatorname{\mathbb{P}}(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{A}}|{\mathcal{F}}_{t})\leq\operatorname{\mathbb{P}}(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t})+\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})]\leq n^{-\min\{\gamma+2\beta-1,\beta-1/2\}+o(1)}.$
[putting everything together]
Lemma 5.4.
Let $v\in V$. Then
$\operatorname{\mathbb{P}}\left(\left||{\mathcal{I}}[T(v),T(v)+n^{-\beta}]|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)])|\right|\geq 2n^{\gamma}\right)\leq\begin{cases}n^{-\min\{\beta,\gamma+2\beta-1\}+o(1)},&\text{$G$ is a tree}.\\ n^{-\min\left\{\beta-\frac{1}{2},\gamma+2\beta-1\right\}+o(1)},&G\in...\end{cases}$
Proof.
Fix $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Define the stopping time $T:=T(v)$, as well as the events
${\mathcal{E}}_{t}$ :=$:=\left\{\left||{\mathcal{I}}[t,t+n^{-\beta}]|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\leq 2n^{\gamma}\right\}$
${\mathcal{E}}$ :=$:=\left\{\left||{\mathcal{I}}[T,T+n^{-\beta}]|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T])|\right|\leq 2n^{\gamma}\right\}.$
Since ${\mathcal{I}}[t,t+n^{-\beta}]={\mathcal{I}}_{1}\cup{\mathcal{I}}_{2}$, Lemmas [ref] imply that if $G$ is a tree,
$\operatorname{\mathbb{P}}({\mathcal{E}}_{t}^{c}|{\mathcal{F}}_{t})\leq n^{-\min\{\beta,\gamma+2\beta-1\}+o(1)}.$
In the general case, we have that
$\operatorname{\mathbb{P}}({\mathcal{E}}_{t}^{c}\cap{\mathcal{A}}|{\mathcal{F}}_{t})\leq n^{-\min\{\beta-1/2,\gamma+2\beta-1\}+o(1)}.$
Since $T$ is almost surely finite, we have that
[…]
if $G$ is a tree. Taking an expectation over ${\mathcal{F}}_{T}$ proves the lemma in the case of trees. Following the same arguments also leads to the claimed result in the general case. ∎
6 Organizing second deriviative stuff
Notation: For $i\in[K]$ and $v\in V$, define
$\widehat{\deg}_{i}(v):=n^{\beta}\left(|{\mathcal{I}}_{i}[T_{i}(v),T_{i}(v)+n^{-\beta}]|-|{\mathcal{I}}_{i}[T_{i}(v)-n^{-\beta},T_{i}(v)]|\right).$
Main result (formulate as algorithm): under some conditions on $\alpha,\beta,K$, it holds with probability $1-o(1)$ that:
  • For $v\in V$ such that $\deg(v)\geq n^{\alpha}$, $\sum_{i=1}^{K}\mathbf{1}(\widehat{\deg}_{i}(v)\geq n^{\alpha})>K/2.$
  • For $v\in V$ such that $\deg(v)\leq d$, $\sum_{i=1}^{K}\mathbf{1}(\widehat{\deg}_{i}(v)\geq n^{\alpha})<K/2.$
[first intermediate result: estimating the cut]
Lemma 6.1.
For any $t\geq 0$, it holds that
($\operatorname{\mathbb{P}}\left(\left\{\left|I\left[t,t+\frac{1}{n^{\beta}}\right]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-c}$
[estimating changes in the cut]
Lemma 6.2.
Let $\gamma>1-\beta$. Then for any $t\geq 0$,
($\operatorname{\mathbb{P}}\left(\left\{\left|\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{2}}\right)\right)-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{\beta}}\right)\right)\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-c}$
[discretizing] Let $\mathbb{T}_{n}$ be a discrete set of evenly spaced points that are $1/n^{2}$ apart.
Lemma 6.3.
Let $K\geq...$. With probability $1-o(1)$, it holds for all $t\in\mathbb{T}_{n}$ that
$\sum_{i=1}^{K}\mathbf{1}\left(\left|I\left[t,t+\frac{1}{n^{\beta}}\right]-I\left[t-\frac{1}{n^{\beta}},t\right]-\frac{\lambda}{n^{\beta}}\left(\operatorname{\mathsf{cut}}({\mathcal{I}}(t))-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{2}}\right)\right)\right)\right|\leq n^{\gamma}\right)\geq\frac{4K}{5}.$
[from here, we need to relate the stopping times to the discretization. The result should then follow immediately.]
[good notation for empirical second derivative?]
[non-tree results below]
[first main result: computing the cut]
Lemma 6.4.
Let $v\in V$ and $\beta>1/2$. It holds that
($\operatorname{\mathbb{P}}\left(\left\{\left|\left|{\mathcal{I}}\left(T(v),T(v)+\frac{1}{n^{\beta}}\right]\right|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
[From here on out, the proof is pretty easy. If you look at $\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)])-\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(u)]$ for consecutively infected vertices $v,u$, you’ll reveal the degree info. After all, if vertices are consecutive, you extract only the effect of the added vertex.]
6.1 Analysis of $|{\mathcal{I}}_{1}|$: Proof of Lemmas 5.1 and 6.4
Our next result shows that with high probability, ${\mathcal{I}}_{2}=\emptyset$.
@@ 6 Organizing second deriviative stuff
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
@@
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Observe that if ${\mathcal{I}}_{2}\neq\emptyset$, then there must exist a path $P$ of length 2 and weight at most $\delta$ for which only the first vertex in $P$ is an element of ${\mathcal{I}}[0,t]$. Additionally, if we define the event
${\mathcal{B}}_{t}:=\{\mathsf{highdeg}(G)\cap{\mathcal{I}}(t,t+\delta]=\emptyset\},$
then all vertices in $P$ must have degree at most $d$. Otherwise, there would exist a path of weight at most $\delta$ connecting a vertex in ${\mathcal{I}}[0,t]$ to some $u\in\mathsf{highdeg}(G)$, which would imply that $T(u)\in(t,t+\delta]$. However, this cannot happen on the event ${\mathcal{B}}_{t}$. Importantly, ${\mathcal{B}}_{t}$ holds with high probability provided ${\mathcal{A}}$ holds; indeed, we have that
@@
To see the implications of the discussion above, let ${\mathcal{Q}}$ be the set of paths $P$ in $G$ of length 2 for which the first vertex in $P$ is the only one in ${\mathcal{I}}[0,t]$, and all vertices in $P$ have degree at most $d$.

### 06 Mar 2024 v2

@@ 3 Preliminaries: Cascade dynamics and infection times
Define the event
${\mathcal{A}}_{t}:=\left\{\text{For all $v\in V\setminus{\mathcal{I}}(t)$, it holds that $|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\leq n^{1/2+o(1)}$}\right\}.$
@@
($\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{J}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)$ ($\leq\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{J}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}_{t}\right|{\mathcal{F}}_{t}\right)$
$=\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{J}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)\mathbf{1}({\mathcal{A}}_{t}).$ (5.3)
We proceed by bounding the probability on the right hand side. For each $v\in\partial{\mathcal{I}}(t)$, notice that
$\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}(t)}F(u,v)\sim\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|),$
@@
$\operatorname{\mathbb{P}}\left(v\in{\mathcal{J}}_{1}|{\mathcal{F}}_{t}\right)=1-\exp\left(-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\right).$
Moreover, by the independence of edge weights, we have that $|{\mathcal{J}}_{1}|\stackrel{{\scriptstyle d}}{{=}}\sum_{v\in\partial{\mathcal{I}}(t)}X(v)$, where the $X(v)$’s are independent with $X(v)\sim\operatorname{Bern}\left(1-e^{-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|}\right)$. We can then bound
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{J}}_{1}|-\operatorname{\mathbb{E}}[|{\mathcal{J}}_{1}||{\mathcal{F}}_{t}]\right|\geq\frac{1}{2}n^{\gamma}\right|{\mathcal{F}}_{t}\right)$ $\stackrel{{\scriptstyle(a)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-\exp(-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|)\right)+n^{\gamma}/6}\right)$
@@
Above, $(a)$ is due to Bernstein’s inequality. Step $(b)$ uses the inequality
$\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-e^{-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|}\right)\leq\frac{\lambda}{n^{\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\leq\frac{\lambda}{n^{\beta}}\left(dn+nm\right)=n^{1-\beta+o(1)},$ (5.5)
where the second inequality above holds since there are at most $n$ vertices $v$ with $|{\mathcal{N}}(v)|\leq d$ and $m$ vertices $v$ with $|{\mathcal{N}}(v)|\leq n$; the final expression uses that $d=n^{o(1)}$ and $m=n^{o(1)}$. Finally, $(c)$ is due to our assumption that $\gamma>(1-\beta)/2$.
It remains to study $\operatorname{\mathbb{E}}[|{\mathcal{J}}_{1}|{\mathcal{F}}_{t}]$. We have that
$\left|\operatorname{\mathbb{E}}[|{\mathcal{J}}_{1}||{\mathcal{F}}_{t}]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|$ $=\left|\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-\exp(\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|)-\frac{\lambda}{n^{\beta}}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\right)\right|$
$\leq\sum_{v\in\partial{\mathcal{I}}(t)}\left|1-\exp(\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|-\frac{\lambda}{n^{\beta}}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\right|$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{\lambda^{2}}{n^{2\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|$
@@
Above, $(c)$ is due to Taylor’s theorem (in particular, the inequality holds as long as $\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|=o(1)$, which is the case since ${\mathcal{A}}_{t}$ holds and $\beta>1/2$). Step $(d)$ follows from the bound $\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\leq n^{1+o(1)}$, which was proved in (5.5). Finally, $(e)$ follows since $\gamma>(1-\beta)/2>1-2\beta$ (where the second strict inequality holds when $\beta>1/2$).
Combining (5.3), (5.4) and (5.6), and then taking an expectation over ${\mathcal{F}}_{t}$ proves the lemma. ∎
Our next result shows that with high probability, ${\mathcal{J}}_{2}=\emptyset$.
@@
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{J}}_{2}|\geq 1\}\cap{\mathcal{A}}\right)\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
@@
As a shorthand, let $\delta:=n^{-\beta}$. Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Observe that if ${\mathcal{J}}_{2}\neq\emptyset$, then there must exist a path $P$ of length 2 and weight at most $\delta$ for which only the first vertex in $P$ is an element of ${\mathcal{I}}(t)$. Additionally, if we define the event
${\mathcal{B}}_{t}:=\{\forall v\in\mathsf{highdeg}(G),T(v)\notin(t,t+\delta]\},$
then all vertices in $P$ must have degree at most $d$. Otherwise, there would exist a path of weight at most $\delta$ connecting a vertex in ${\mathcal{I}}(t)$ to some $u\in\mathsf{highdeg}(G)$, which would imply that $T(u)\in(t,t+\delta]$. However, this cannot happen on the event ${\mathcal{B}}_{t}$. Importantly, ${\mathcal{B}}_{t}$ holds with high probability provided ${\mathcal{A}}$ holds; indeed, by Lemma 3.8 we have that
@@
To see the implications of the discussion above, let ${\mathcal{Q}}$ be the set of paths $P$ in $G$ of length 2 for which the first vertex in $P$ is the only one in ${\mathcal{I}}(t)$, and all vertices in $P$ have degree at most $d$.